OpenStudy (anonymous):

How many solutions does this system have? -3x+6y=10 -3x+6y=-4

OpenStudy (anonymous):

the left hand sides are the same the right hand sides are different

OpenStudy (anonymous):

Here are some hints. If you solve it and you get: ~0y + 0x =0 , this means it will have an infinite solution ~0y +0x = a, where a is any number, this means that it will not have any solutions because 0 cannot equal to "a". ~if there's a value for x and y, then it will have one solution... i hope i'm right (: so what did you get when you solve this system?

OpenStudy (anonymous):

there is no way for \(-3x+6y\) to be two different numbers with the same x and y

OpenStudy (anonymous):

just like you can't have both \(2x=5\) and \(2x=7\) for the same \(x\)

OpenStudy (anonymous):

i honestly dont even know how to solve this lol

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

don't worry we're here to help! there are two methods of solving systems of equation: \(\sf substitution\), where you have to evaluate one equation in terms of one variable and plug it in in the second equation. \(\sf elimination\) where you have to eliminate one variable to solve for the other. now with your problem, elimination method is the best way. -3x+6y=10 -3x+6y=-4 what do you have to do to eliminate one variable when you'll add the two equations? since they already look 'similar', no need to do anything, you just have to simply add the two equations, can you do that?

OpenStudy (anonymous):

what do you mean by simplify?

OpenStudy (anonymous):

oh nevermind

OpenStudy (anonymous):

i read it wrong lol hold on

OpenStudy (anonymous):

**my bad.. i made a mistake.. equations are similar.. in order to eliminate one variable, you have to multiply one equation by -1... let's say eqan1: -1(-3x+6y=10 ) so you'll have 3x-6y =10 now you have : 3x-6y =-10 -->eqn 1 -3x+6y=-4 --> eqn 2 you need to do : eqn 1 + eqn 2